### v1

# Gait Transitions in Load-Pulling Quadrupeds: Insights from Sled Dogs and a Minimal SLIP Model

Abstract
Quadrupedal animals employ diverse galloping strategies to optimize speed, stability, and energy efficiency. However, the biomechanical mechanisms that enable adaptive gait transitions during high-speed locomotion under load remain poorly understood. In this study, we present new empirical and modeling insights into the biomechanics of load-pulling quadrupeds, using sprint sled dogs as a model system. High-speed video and force recordings reveal that sled dogs often switch between rotary and transverse galloping gaits within just a few strides and without any observable changes in speed, stride duration, or terrain, providing clear evidence of locomotor multistability during high-speed load-pulling. To investigate the mechanical basis of these transitions, a physics-based quadrupedal Spring-Loaded Inverted Pendulum model with hybrid dynamics and prescribed footfall sequences to reproduce the asymmetric galloping patterns observed in racing sled dogs. Through trajectory optimization, we replicate experimentally observed gait sequences and identify swing-leg stiffness modulation as a key control mechanism for inducing transitions. This work provides a much-needed biomechanical perspective on high-speed animal draft and establishes a modeling framework for studying locomotion in pulling quadrupeds, with implications for both biological understanding and the design of adaptive legged systems.
DOI
TBD.
Keywords
Gait transitions in quadrupeds,   Load-pulling locomotion dynamics,   Sled dogs,
Spring-loaded inverted pendulum modeling
preprint: PRX LIFE
I Introduction
Figure 1: Sled dog gait analysis. (A) A video frame showing sled dogs in motion, with colored markers used to indicate each foot. Similar markers were used to identify stance phases during the stride, labeled manually. (B) A segment of reconstructed footfall data and recorded tugline force. The stance-stride plot illustrates footfall timing, with the colored bars representing the stance durations for each limb. Cartoons depict the ordering of the legs in the corresponding rotary and transverse gait patterns present. The synchronized force recording measures the dynamic loading of the tugline throughout the sled dogs’ gait.
Coordinated limb movement is fundamental to effective animal locomotion and has long inspired the design of agile, efficient robotic systems [1, 2]. Among terrestrial mammals, quadrupedal galloping is one of the most mechanically dynamic and energetically demanding gait patterns. Species such as dogs, horses, and cheetahs adopt distinct galloping strategies—particularly rotary and transverse variants—to optimize performance across different speeds and terrains [3, 4]. Traditionally viewed as a stereotyped, high-speed gait with transitions linked to velocity or environmental changes [5], galloping has more recently been reinterpreted through the lens of locomotor multistability, where multiple gait solutions can coexist under nearly identical conditions [6]. For instance, modeling studies have shown that even small variations in limb compliance or body symmetry can give rise to distinct galloping modes within the same speed range [7, 8]. However, such phenomena have rarely been examined in quadrupeds operating under external load. Sled dogs offer a particularly compelling and under-studied example of this phenomenon: they sustain high-speed galloping while pulling a sled, requiring coordination not only for self-propulsion but also to generate forward force along a connecting rope known as the tugline. This added load introduces unique biomechanical constraints, as the resulting tension likely affects stance duration, limb loading, and possibly interlimb timing.
Despite growing interest in quadrupedal locomotion, particularly in working contexts, the dynamics of load-pulling quadrupeds remain poorly understood. A long history of animal draft in agriculture, transportation, and sport has generated limited biomechanical literature on load-pulling quadrupeds, especially under dynamic conditions. Most existing work has focused on practical questions in agricultural settings, emphasizing work output or harness efficiency during heavy-load, static-pulling scenarios [9, 10, 11, 12, 13]. Recent robotics research has begun to model dynamic quadruped-load systems [14], but such models typically operate at the path-planning level and do not capture limb-level dynamics or the energetic coupling between the quadruped and its load. To date, no dynamical model has been developed that captures the limb-load coupling or stride-level coordination necessary to replicate the complex dynamics of quadruped-load systems, especially those involving gait transitions.
In this study, we contribute new empirical and modeling insights to the study of load-pulling quadrupeds. Using high-speed video recordings of racing sled dogs, we reconstructed footfall sequences and tugline forces during high-speed locomotion. Our analysis reveals that sprint sled dogs do not rely on a single fixed gallop. Instead, they frequently switch between rotary and transverse galloping patterns, often within a few strides and without apparent external triggers, demonstrating a high degree of locomotor multistability.
We build on the Spring-Loaded Inverted Pendulum (SLIP) framework [15, 16, 17], a reduced-order model widely used to study locomotor dynamics, and extend it to quadrupedal systems that incorporate limb asymmetry, external load pulling, and predefined footfall sequences. While prior models have explored gait energetics [18] and bipedal transitions [19], our formulation captures the full-body dynamics and interaction forces between a quadruped and a load. In contrast to neuromechanical simulations that require detailed anatomical modeling [20], our model achieves generalizability with fewer parameters while remaining grounded in physical dynamics. To our knowledge, this is the first dynamical model capable of reproducing not only the detailed stride mechanics but also the spontaneous gait transitions observed in load-bearing quadrupeds. By embedding empirical gait sequences and optimizing against experimental data, our model provides new insight into the control strategies and mechanical feasibility of stride-to-stride transitions in pulling contexts.
II Gait Transitions in Sprint Sled Dogs
II.1 Gait Pattern Identification
Fieldwork was conducted in collaboration with a professional sprint-racing sled dog kennel, with data collection efforts integrated into their regular training routine. Sprint-racing is a fast-paced style of dogsled racing, in which dogs typically maintain galloping speeds of 25 to 40 km per hour over distances less than 50 km [21]. Training runs took place on dirt trails or shallow snow, with teams pulling a wheeled training rig carrying one to two human passengers.
Precise measurement of limb-specific stance and stride timing typically relies on controlled experiments using high-speed video, with either manual annotation [22, 23] or automated labeling methods [24]. However, replicating such methods in the field, especially with free-running sled dogs, is logistically challenging. To overcome this, we mounted an action camera (GoPro, 60 fps +) on an extended pole held from the training rig. This setup enabled us to capture targeted side-view footage of two individuals—referred to here as Individual 1 and Individual 2—during select training runs. From this footage, we manually annotated stance and stride durations for each limb (Fig. 1A).
These annotations were used to classify gait patterns based on the unique sequence of limb-ground interactions. Following Hildebrand’s convention [4], a footfall diagram can be used to depict contact durations as solid bars and aerial phases as gaps. A short selection of footfalls, presented this way (Fig. 1B), illustrates the frequent switching between transverse and rotary gallops during high-speed pulling. In transverse gallop, the hind legs strike in sequence, followed by the contralateral and then ipsilateral forelegs. In rotary gallop, the placement of the second hind foot is followed by the ipsilateral forefoot, creating a seemingly rotating sequence of footfalls around the body [3]. All four permutations of these galloping patterns are illustrated in Fig. 2.
Our analysis also benefits from data collected in parallel with video using custom data loggers integrated into the dogs’ harnesses. These sensors were clipped onto a standard ‘X-back’ harness near the dog’s withers, and measured tugline force (62 Hz) and tri-axial acceleration (125 Hz) throughout each stride. After experiments were completed, data was transferred from the data loggers to a computer for post-hoc analysis. For this paper, only the force data was analyzed. For a more detailed overview of the sensors used, including design files and firmware, see  [^1] https://github.com/DLARlab/2025 Gait Transitions in Load Pulling Quadrupeds Insights.git. The GoPro camera recorded GPS data in the video file metadata, and the data logger timestamped samples using an included GPS module. The shared epoch timestamps allowed force measurements to be synchronized with video recordings. Fig. 1B shows this tension over a short period of time. Notably, the measured force exhibits stride-synchronized oscillations, with peaks that often—but not always—coincide with the touch-down of the trailing forelimb. Load cells were calibrated between runs using fixed weights to correct for mechanical drift or deformation.
II.2 Gait Pattern Definitions
Figure 2: Quadrupedal footfall patterns following Hildebrand’s convention [4]. The horizontal axis represents normalized stride time. Colored bars denote ground contact phases for each leg: green (LH), red (LF), yellow (RF), and blue (RH). Each row represents a distinct galloping gait, with white backgrounds for transverse sequences and grey for rotary. Touchdown ($t_{i}^{\scriptstyle\text{TD}}$) and liftoff ($t_{i}^{\scriptstyle\text{LO}}$) define the stance period. Phase lags are annotated for fore-hind ($d_{p}$), hindlimb ($d_{h}$), and forelimb ($d_{f}$) timing. Below each bar plot, cartoons depict limb configurations, and adjacent dot plots show footfall sequences, with green and blue font labels for left- and right-leading gaits.
Each galloping gait is defined by the touchdown (TD) and liftoff (LO) events of all four limbs over a single stride cycle. We assume that each leg undergoes exactly one stance and one flight phase per stride. Limb labels are assigned as follows: left hind (LH), left fore (LF), right fore (RF), and right hind (RH).
To standardize comparisons across gait types, we define an extended flight phase between the LO of the trailing hindlimb and the TD of the leading forelimb, during which the torso is airborne. The start of each stride is anchored to the midpoint of this aerial phase, corresponding to the moment of peak torso height. This convention follows biomechanical interpretations of galloping dynamics and ensures a consistent phase reference for stride analysis [7, 26]. Each stride is encoded by the timing vector:
[…]
Each element $t_{i}$, for $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$, is normalized to the stride interval $[0,1)$ w.r.t. the total stride time $T$ (in seconds). The timing vector is organized by anatomical order rather than temporal sequence, ensuring consistent limb labeling across gait types.
Using the timing vector, we define several normalized metrics to quantify key temporal relationships:
  • Duty factor: $d_{t}\in[0,1)$ is the stance duration for a given leg:
    $d_{t}:=(t^{\scriptstyle\text{LO}}_{i}-t^{\scriptstyle\text{TD}}_{i})\bmod 1.$
    The modulo handles wrapping when liftoff occurs before touchdown in normalized stride time.
  • Forelimb phase lag: $d_{f}\in(-0.5,0.5]$ measures the TD timing difference between left and right forelimbs:
    $d_{f}:=\left[\left(t^{\scriptstyle\text{TD}}_{\scriptstyle\text{LF}}-t^{\scriptstyle\text{TD}}_{\scriptstyle\text{RF}}+0.5\right)\bmod 1\right]-0.5.$
  • Hindlimb phase lag: $d_{h}\in(-0.5,0.5]$ is defined similarly:
    $d_{h}:=\left[\left(t^{\scriptstyle\text{TD}}_{\text{LH}}-t^{\scriptstyle\text{TD}}_{\scriptstyle\text{RH}}+0.5\right)\bmod 1\right]-0.5.$
  • Fore-hind phase lag: $d_{p}\in[-0.5,0.5)$ compares average mid-stance timing between hindlimbs and forelimbs:
    $d_{p}:=\left[\frac{1}{2}\left(\bar{t}_{\scriptstyle\text{LH}}+\bar{t}_{\scriptstyle\text{RH}}-\bar{t}_{\scriptstyle\text{LF}}-\bar{t}_{\scriptstyle\text{RF}}\right)+0.5\right]\bmod 1-0.5,$
    where $\bar{t}_{i}:=\frac{1}{2}(t^{\scriptstyle\text{TD}}_{i}+t^{\scriptstyle\text{LO}}_{i})$ denotes the mid-stance time of leg $i$.
With the timing variables and phase lag metrics defined above, we now define the space of asymmetric galloping gaits as those in which both forelimb and hindlimb phase lags are neither zero nor equal to $0.5$:
$G_{e}=\left\{\bm{t}\,\middle|\,d_{f},\ d_{h}\notin\{0,0.5\}\right\}.$ (2)
Phase lags of zero correspond to bounding or half-bounding gaits, while values of $0.5$ indicate symmetrical gaits such as pacing or trotting, following Hildebrand’s classification [4]. Within this space, each canonical galloping gait $G_{e}$, for $e\in\{\textit{TL},\textit{TR},\textit{RL},\textit{RR}\}$, is uniquely determined by the sign pair $(d_{f},d_{h})$:
  • TL (transverse, left-leading): $d_{f}>0,\ d_{h}>0$
  • TR (transverse, right-leading): $d_{f}<0,\ d_{h}<0$
  • RL (rotary, left-leading): $d_{f}>0,\ d_{h}<0$
  • RR (rotary, right-leading): $d_{f}<0,\ d_{h}>0$
These gait labels are visually depicted in Fig. 2. Transverse gallops appear in the top row (white background), and rotary gallops in the bottom row (grey background). Colored bars indicate stance phases; schematic illustrations below each plot highlight limb movement; and dot plots emphasize the sequence of footfalls and leading limb. Gait transitions correspond to sign changes in either $d_{f}$ or $d_{h}$. For instance, a transition from TL to RL involves a reversal in hindlimb phasing (change in $d_{h}$), while a switch from RL to RR reflects a forelimb phase reversal (change in $d_{f}$). Up to 12 distinct transitions can be theoretically defined among the four canonical galloping gaits based on the sign combinations of phase lags. These transitions form the basis for analyzing discrete gait-switching behaviors observed in sled dogs.
III Hybrid SLIP-Based Modeling of Load-Pulling Galloping Gaits
This section presents a physics-based modeling framework designed to reproduce and analyze galloping gait patterns in sled dogs pulling a load. We begin by introducing a planar quadrupedal SLIP model extended with a tugline-coupled load, followed by a hybrid dynamics formulation that captures discrete gait transitions through leg touchdown and liftoff events. We then describe an optimization-based approach for identifying periodic solutions that closely match experimentally observed footfall timings and tugline forces.
III.1 Quadrupedal SLIP Model with Load
We extend a planar SLIP model to capture the dynamics of sled dogs pulling a load at high speed. The base model, adapted from our previous work [27, 7], consists of a rigid torso supported by four massless spring legs and augmented with a discrete load mass connected via a unilateral spring element that mimics the intermittent tension of a tugline.
Figure 3: The quadrupedal SLIP model used in this work, with legs connected to the torso through both linear and torsional springs. The model is constrained to planar sagittal motion. Each leg is passively actuated and exhibits distinct swing and stance dynamics. The pulling load is attached via a unilateral spring element modeling the tugline.
The quadruped consists of a rigid body of mass $M_{\text{torso}}$ and moment of inertia $J_{\text{torso}}$, supported by four limbs anchored at fixed hip and shoulder joints. The system evolves in the sagittal plane and is described by eight generalized coordinates: the torso position $(q_{x},q_{z})$, pitch angle $q_{\theta}$, four leg swing angles $q_{\alpha,i}$ for $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$, and the relative horizontal displacement of the load $q_{\text{load}}$.
The model forms a hybrid dynamical system, with stance and flight phases governed by distinct sets of equations. The hind and fore legs are mounted at distances $l_{b,H}$ and $l_{b,F}$ from the torso Center of Mass (CoM), respectively. In stance, ground contact imposes holonomic constraints and generates ground reaction forces (GRFs); in flight, limbs follow unconstrained passive dynamics driven by torsional stiffness. Each leg behaves as a linear spring with stiffness $k_{\text{leg}}$ during stance, storing and releasing energy to propel the body forward. Legs are massless but terminate in point feet with small masses $m_{\text{foot}}$ to define discrete ground contact events [28]. A fore–hind stiffness ratio $r_{\text{leg}}$ is included to model anatomical asymmetry between limb pairs. During flight, each leg swings passively under a torsional spring of stiffness $k_{\text{swing}}$, resulting in approximately sinusoidal motion.
To represent the external load, we introduce a point mass $M_{\text{load}}$ connected to the torso by a unilateral, massless spring (the tugline) with stiffness $k_{\text{tug}}$ and resting length $l_{t}$. The tugline exerts force only when stretched beyond its rest length, producing discontinuous coupling between the quadruped and the load. When slack, the load follows its inertial trajectory; when taut, it exerts a horizontal restoring force that influences both the load and torso dynamics. To maintain steady galloping in the presence of energy loss (e.g., due to sliding friction at the load), similar to the Inverted Pendulum (IP) model with rigid legs [29], we tilt the terrain by a small slope angle $\theta_{\text{slope}}$ and model Coulomb friction at the load-ground interface with coefficient $\mu_{\text{ground}}$. This slope enables gravitational work to balance energy dissipation, facilitating sustained periodic motion. All parameters are summarized in Table 1 and are normalized by the total system mass $m$, gravitational acceleration $g$, and resting leg length $l_{0}$.
Table 1: Key model parameters used in the quadrupedal SLIP model with load. All quantities are normalized using total mass $m$, gravity $g$, and resting leg length $l_{0}$ [28].
Symbol Description Units
$M_{\rm torso}$ Torso mass $m$
$m_{\rm foot}$ Mass of each foot $m$
$J_{\rm torso}$ Torso pitching inertia $ml_{0}^{2}$
$k_{\rm leg}$ Leg spring stiffness $mg/l_{0}$
$r_{\rm leg}$ Stiffness ratio between hind/fore legs $\cdot$
$k_{\mathrm{swing},i}^{n}$ Torsional stiffness of swing‐leg spring $mgl_{0}$
$l_{l}$ Resting leg length $l_{0}$
$l_{b,H}$, $l_{b,F}$ Distance from CoM to hip/shoulder $l_{0}$
$M_{\rm load}$ Load mass $m$
$k_{\rm tug}$ Tugline (unilateral spring) stiffness $mg/l_{0}$
$l_{t}$ Tugline resting length $l_{0}$
$\mu_{\text{ground}}$ Friction coefficient of ground contact $\cdot$
$\theta_{\text{slope}}$ Slope angle of running surface $\cdot$
To enable gait transitions and capture limb-specific asymmetries observed in natural locomotion, each leg is assigned an independent swing-leg stiffness parameter $k_{\rm swing}$, which may vary between consecutive strides. For an $N$-stride sequence, the stiffness set is denoted as $\bm{k}_{\rm swing}=\{k_{\mathrm{swing},i}^{n}\}$, where $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$ and $n=1{:}N$. This formulation supports stride-wise modulation of limb stiffness and enables replication of experimentally observed gait transitions driven by localized changes in limb dynamics.
III.2 Hybrid Dynamics of the Quad-load System
The quadrupedal SLIP model with tugline evolves as a hybrid dynamical system, where continuous dynamics change with discrete contact states. The configuration vector is defined as $\bm{q}=[q_{x},\ q_{z},\ q_{\theta},\ \bm{q}_{\alpha}^{\top},\ q_{\text{load}}]^{\top}\in\mathbb{R}^{8},$ and the continuous dynamics are governed by the Euler-Lagrange equations:
$\bm{M}(\bm{q})\ddot{\bm{q}}+\bm{C}(\bm{q},\dot{\bm{q}})\dot{\bm{q}}+\bm{G}(\bm{q})=\sum_{i}\bm{J}_{i}^{\top}(\bm{q})\bm{\lambda}_{i},$ (3)
where $\bm{M}(\bm{q})\in\mathbb{R}^{8\times 8}$ is the mass-inertia matrix, $\bm{C}(\bm{q},\dot{\bm{q}})\dot{\bm{q}}$ captures Coriolis and centrifugal terms, and $\bm{G}(\bm{q})$ encodes gravitational forces. Each contact force $\bm{\lambda}_{i}\in\mathbb{R}^{2}$ is transmitted through a Jacobian $\bm{J}_{i}(\bm{q})\in\mathbb{R}^{2\times 8}$, which maps GRFs into generalized coordinates.
We assume that each leg undergoes exactly one touchdown and one liftoff per stride to avoid Zeno behavior [30, 31]. As a result, the eight timing variables $t_{i}$ in (1) partition the hybrid gait cycle into domains $j=1{:}9$, each corresponding to a unique combination of ground-contacting legs. Within each domain, the associated holonomic constraints are enforced using stacked Jacobians $\bm{J}_{j}$ [18], yielding the following differential-algebraic formulation:
$\mathcal{F}_{j}:\begin{bmatrix}\bm{M}(\bm{q})&-\bm{J}_{j}^{\intercal}(\bm{q})\\ \bm{J}_{j}(\bm{q})&\bm{0}\end{bmatrix}\begin{bmatrix}\ddot{\bm{q}}\\ \bm{\lambda}_{j}\end{bmatrix}=\begin{bmatrix}-\bm{C}(\bm{q},\dot{\bm{q}})\dot{\bm{q}}-\bm{G}(\bm{q})\\ -\dot{\bm{J}}_{j}(\bm{q},\dot{\bm{q}})\dot{\bm{q}}\end{bmatrix},$ (4)
where $\bm{\lambda}_{j}$ includes all GRFs in domain $\mathcal{F}_{j}$. Each of the $162^{4}=$ 16 possible contact combinations defines a continuous domain in the hybrid system.
Touchdown and liftoff events for each leg $i$ define the switching surfaces:
$\begin{split}\mathcal{C}^{\scriptstyle\text{TD}}_{i}&=\left\{(\bm{q},\dot{\bm{q}})\in\mathcal{TQ}\,\middle|\,l_{i}(\bm{q})=l_{l},\ \dot{l}_{i}(\bm{q})<0\right\},\\ \mathcal{C}^{\scriptstyle\text{LO}}_{i}&=\left\{(\bm{q},\dot{\bm{q}})\in\mathcal{TQ}\,\middle|\,l_{i}(\bm{q})=l_{l},\ \dot{l}_{i}(\bm{q})>0\right\},\end{split}$ (5)
where $l_{i}(\bm{q})$ is the actual leg length and $l_{l}$ is the rest length. The hybrid model is written as:
:$\mathcal{H}:\left\{\begin{array}[]{ll}\mathcal{F}_{j},&(\bm{q},\dot{\bm{q}})\notin\mathcal{C}_{j}(t),\\ \dot{\bm{q}}^{+}=\Delta_{\mathcal{F}_{j}\rightarrow\mathcal{F}_{j+1}}\dot{\bm{q}}^{-},&(\bm{q},\dot{\bm{q}})\in\mathcal{C}_{j}(t),\end{array}\right.$ (6)
where $\Delta$ denotes the impact map that instantaneously resets velocities at transitions. Further implementation details are provided in [7, 32, 18]
III.3 Trajectory Optimization for Gait Replication
To reproduce periodic gaits and gait transitions observed in sled dogs, we formulate a trajectory optimization problem that identifies feasible motion trajectories of the hybrid SLIP-load model while minimizing a composite cost. The objective function integrates physical consistency, alignment with experimentally recorded footfall timings and tugline forces, and stride duration fidelity.
The optimization is performed over a reduced set of decision variables, including the initial state $(\bm{q}_{0},\dot{\bm{q}}_{0})$ and model parameters $\bm{p}$, which consist of all entries listed in Table 1. Among these, only the torso center-of-mass locations $l_{b,H}$ and $l_{b,F}$ are held fixed to preserve anatomical consistency. Let $N$ denote the number of strides and the objective is expressed as:
$\min_{\begin{subarray}{c}\bm{q}_{0},\ \dot{\bm{q}}_{0},\ \bm{p}\\ \end{subarray}}$ $\ \ \sum_{n=1}^{N}\Big{[}w_{d}\left\|T_{\mathrm{sim}}^{n}-T_{\mathrm{exp}}^{n}\right\|^{2}+w_{t}\left\|\bm{t}_{\mathrm{sim}}^{n}-\bm{t}_{\mathrm{exp}}^{n}\right\|^{2}$
$\ \ +w_{f}\left\|\bm{F}_{\mathrm{sim}}^{n}-\bm{F}_{\mathrm{exp}}^{n}\right\|^{2}+w_{r}\left\|\bm{R}^{n}(\bm{q})\right\|^{2}\Big{]},$ (7)
where $\bm{t}_{\mathrm{sim}}^{n}$, $\bm{F}_{\mathrm{sim}}^{n}$, and $T_{\mathrm{sim}}^{n}$ denote the simulated footfall timing vector, tugline force trajectory, and stride duration for stride $n$, respectively. Experimental data are denoted by the corresponding $\mathrm{exp}$ subscripts. The residual vector $\bm{R}^{n}(\bm{q})$ captures physical consistency constraints, including both (i) holonomic constraints enforcing correct leg lengths at touchdown and liftoff events, and (ii) continuity constraints for state variables across stride boundaries. The weights $w_{r},w_{t},w_{f},w_{d}$ are manually tuned to balance the influence of each term.
Single-Stride Optimization: When $N=1$, the optimization produces a periodic gait by enforcing cycle closure constraints:
$\bm{q}(t_{0})=\bm{q}(t_{\mathrm{end}}),\quad\dot{\bm{q}}(t_{0})=\dot{\bm{q}}(t_{\mathrm{end}}),$
ensuring that the simulated trajectory forms a dynamically consistent limit cycle matching experimental stance patterns and tugline force profiles.
Multi-Stride Optimization: For gait transitions and sequence replication, the trajectory is segmented into $N$ strides indexed by $n=1{:}N$, with each stride spanning $t_{0}^{n}$ to $t_{\mathrm{end}}^{n}$. Continuity is enforced between successive strides:
$\bm{q}(t_{0}^{n+1})=\bm{q}(t_{\mathrm{end}}^{n}),\quad\dot{\bm{q}}(t_{0}^{n+1})=\dot{\bm{q}}(t_{\mathrm{end}}^{n}).$
To replicate experimentally observed stride-to-stride gait changes, the swing stiffness parameters $k_{\mathrm{swing},i}^{n}$ for $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$ are treated as stride-specific optimization variables and allowed to vary across strides. Initial guesses are manually constructed based on approximate periodic solutions and refined using a direct collocation method combined with nonlinear programming. This procedure yields dynamically feasible trajectories that closely reproduce experimental stride patterns and tugline force behaviors for both steady-state and transitioning gaits.
IV Results
In this section, we present the analysis of general patterns observed in galloping gaits used by load dogs during load pulling. We observed that load dogs frequently and randomly transition among four different types of galloping gaits. To reveal the underlying mechanism behind these frequent and random gait transitions, we applied our quad-load model optimization approach to achieve two primary objectives. First, we aimed to replicate the quad-load motion accurately for a single stride, demonstrating our model’s capability to reproduce realistic locomotion dynamics. Second, we conducted simulations replicating the quad-load motion across different galloping gait types. In this second part, selected model parameters were allowed to vary freely between strides, enabling us to uncover the mechanisms driving gait transitions by analyzing parameter variations.
IV.1 Observed Distribution of Gait Patterns
Our analysis of high-speed video recordings revealed that sled dogs frequently switch between different galloping gaits during load-pulling. Fig. 4 summarizes the observed distribution of gallops and the frequency of transitions between them for two individual dogs. The arrangement of gait types in the figure follows the same convention as Fig. 2, with transverse gaits (TL, TR) placed above and rotary gaits (RL, RR) below, and left- and right-leading gaits positioned to the left and right, respectively. The size of each dot is proportional to the number of strides recorded for that gait, while the thickness of the connecting arrows reflects the frequency of transitions between gaits.
Data for Individual 1, compiled from 189 recorded strides, reveals the use of all four galloping gait patterns. Among these, RL was the most frequently used, occurring in 120 out of 190 strides. This was followed by TR, which appeared in 50 strides. In contrast, TL and RR were used only occasionally. Gait transitions were unevenly distributed. For RL, 79.83% of consecutive strides remained in the same gait. Transitions to TR occurred in 15.13% of cases, and to TL in 5.04%. For strides beginning in TR, 44.00% remained in TR, 50.00% transitioned to RL, and only 6.00% transitioned to TL. No transitions to RR were observed from TR either. These statistics suggest a strong preference for RL and TR, while TL and RR appear to be infrequent or short-lived. The most common transitions occurred between RL and TR (RL $\leftrightarrow$ TR), which require changing the touchdown order of only the hindlimbs. In contrast, transitions between RL and RR were never observed, likely because they involve a simultaneous reordering of both the forelimb and hindlimb touchdown sequences within a single stride. This more complex coordination change may be less favorable during steady high-speed locomotion.
Data for Individual 2, compiled from 195 recorded strides, reveals a highly consistent gait preference. Nearly all strides were executed in the TR gait, with only a small fraction performed in RL. Specifically, TR accounted for 209 out of 211 strides, while RL was observed in just a single stride. The other two gaits, TL and RR, were not used at all in the recorded dataset. No transitions were observed among three of the four gait types, as TR and RL were the only patterns present. For strides beginning in TR, 99.52% remained in the same gait, while only 0.48%, corresponding to a single stride, transitioned to RL. Since RL occurred only once and was not followed by any additional strides of that type, no outbound transitions from RL could be evaluated. This consistency indicates a strong and stable preference for the TR gait in Individual 2, with other patterns either avoided or unsustainable. The lack of bidirectional transitions suggests that, in contrast to Individual 1, this dog maintained a highly stable coordination strategy with minimal gait variability across the recorded session.
Figure 4: Distribution and transition dynamics of galloping gaits used by Individual 1 (A) and Individual 2 (B). Each node represents a distinct gait type, with the dot size proportional to its overall usage frequency. Directed arrows indicate transitions between gaits, with arrow thickness representing transition frequency. For each gait node, a corresponding dot plot is shown to illustrate the footfall sequence, using the same color (green for left-leading, blue for right-leading) and background (white for transverse, grey for rotary) convention as defined in Fig. 2. This integrated view links gait prevalence, transitions, and sequence structure in a unified representation.
IV.2 Single Stride Replication
In this subsection, we applied the optimization algorithm described in Section III.3 to replicate the four galloping gaits observed in our dataset. We compared footfall timings and tugline forces for both Individual 1 and Individual 2 across the gaits RL, RR, TL, and TR.
For Individual 1, the two most frequently used gaits, RL and TR, are shown in Fig. 5 for detailed comparison between empirical observations and model simulation. Mean touchdown and liftoff timings were extracted from the recorded stride sequences. The averaged stance durations for each leg are shown as white bars in the left panels of Fig. 5, with error bars indicating variability. Simulated stance durations are overlaid as colored bars for direct comparison. In the right panels, tugline forces are shown as solid blue curves for the simulation and as dashed red curves with shaded bands indicating the measured mean and variance.
Figure 5: Single-stride solutions for Individual 1, comparing two gaits: (A) RL and (B) TR. Left subplots show experimental and simulated footfall sequences; right subplots illustrate experimental and simulated tugline forces along the tugline. Experimental means (solid curves) and variances (shaded regions and error bars) are computed from 120 (RL) and 50 (TR) samples.
Table 2: Summary of $R^{2}$ evaluations for footfall timing ($R^{2}_{t}$), tugline force ($R^{2}_{F}$), and stride duration ($R^{2}_{D}$) across different gait types and individuals. The number of strides used to compute each experimental average is listed in the rightmost column. Data for RL and TR for Individual 1 correspond to those shown in Fig. 5.
Dog Name Gait Type $R^{2}_{t}$ $R^{2}_{F}$ $R^{2}_{D}$ Strides
Individual 1 TL 0.99 0.95 0.93 9
Individual 1 TR 0.99 0.96 0.99 50
Individual 1 RL 0.99 0.96 1.00 120
Individual 1 RR 0.99 0.95 0.95 10
Individual 2 TR 0.99 0.95 0.92 209
Individual 2 RL 0.99 0.96 0.99 2
For the RL gait of Individual 1 (Fig. 5A), the simulated stance onset and offset closely match the experimental profiles across all limbs. The mean stance duration error is below 3% of the stride cycle for each leg. The tugline force profile also aligns well, with the simulated curve capturing both the timing and amplitude of the force peaks. The peak force in simulation reaches approximately $0.27\,mg$, consistent with the experimental average. Minor discrepancies appear in the force rising phase but remain within experimental variability. This high degree of agreement is reflected in Table 2, where RL yields $R^{2}=0.99$ for footfall timing, $R^{2}=0.96$ for tugline force, and a perfect $R^{2}=1.00$ for stride duration.
As for the TR gait of Individual 1 (Fig. 5B), the model again reproduces stance durations with high accuracy. Some asymmetry in the forelimb stance timing is evident in the experimental data and is qualitatively captured by the simulation. Tugline force agreement is similarly strong, with an overall $R^{2}=0.96$ and peak force amplitude closely matching the measured mean. The stride duration agreement is slightly lower than for RL but still excellent, at $R^{2}=0.99$.
Due to space constraints, the optimization results for the remaining gaits and for Individual 2 are not visualized here but are included in the publicly available source code and data repository . Table 2 summarizes the $R^{2}$ values across all gait and individual combinations. Footfall timing predictions are highly consistent, with all combinations achieving $R^{2}_{t}=0.99$, indicating precise alignment of simulated and experimental touchdown and liftoff events. Tugline force predictions also show strong agreement, with $R^{2}_{F}$ values ranging from 0.95 to 0.96 across gaits. Stride duration tracking is most accurate for the frequently used RL and TR gaits, where $R^{2}_{D}$ exceeds 0.99, and is slightly lower for TL and RR, which exhibit more variability. Although TL and RR are less frequently used gaits, the SLIP model still maintains reasonable fidelity, with timing and force $R^{2}$ values above 0.93 in all cases. Slightly reduced agreement in TL, especially for Individual 1, may reflect increased variability in experimental execution or decreased model robustness for rarely used gaits. Nevertheless, the model performs reliably across all conditions tested.
Figure 6: Replicated gait transition for Individual 1 from transverse right (TR) to rotary left (RL) gallop. (A) Simulated footfall sequence. (B) Swing leg angular velocities. (C) Tugline force trajectory. (D) Absolute cost variation under ±$\pm 5\%$ perturbations in $\Delta k_{\text{swing},i}$. (E) Relative percentage cost variation. These results highlight the dominant role of hindlimb stiffness modulation in driving the transition.
IV.3 Gait Transition Replication
[…]
Table 3: Swing stiffness changes ($\Delta k_{\mathrm{swing},\ i}:=k_{\mathrm{swing},\ i}^{n+1}-k_{\mathrm{swing},\ i}^{n}$ $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$) across transitions. Highlighted cells indicate the leg pair responsible for the transition.
Transition Switched $R^{2}_{\text{w}}$ LH RH LF RF
[…]
TR $\leftrightarrow$ RL Gait Transition
As shown in Fig. 6(A), the simulation accurately reproduces the key mechanical features of the TR $\rightarrow$ RL transition, achieving a weighted coefficient of determination of $R^{2}_{\text{w}}=0.95$. Both the footfall sequence and the tugline force trajectory closely match experimental observations across the two-stride transition. The middle panel illustrates swing leg angular velocities, which exhibit near-sinusoidal trajectories during flight—consistent with passive leg oscillations. Since the oscillation period inversely reflects swing stiffness, shorter cycles indicate higher values of $k_{\text{swing}}$.
A clear reversal in hindlimb asymmetry is observed across the transition. In the first stride, the left hind (LH, green) shows a shorter period than the right hind (RH, blue), suggesting higher stiffness. In the second stride, this pattern reverses, with RH now oscillating faster than LH. The forelimb trajectories (LF: red, RF: yellow), by contrast, remain symmetric between strides, indicating stable stiffness. These observations suggest that the transition is primarily driven by differential modulation of hindlimb swing stiffness.
Quantitative analysis supports this interpretation. Defining the stride-to-stride stiffness change as $\Delta k_{\mathrm{swing},\ i}:=k_{\mathrm{swing},\ i}^{n+1}-k_{\mathrm{swing},\ i}^{n}$, we observe substantial and opposing variations in the hindlimbs: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LH}}=-23.8\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RH}}=+23.1\,mgl_{0}$. In contrast, the forelimbs exhibit only minor adjustments: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LF}}=-1.38\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RF}}=+1.57\,mgl_{0}$. These values align with the observed swing leg velocity asymmetries and are summarized in Table 3, where the dominant hindlimb contributions are highlighted.
To evaluate whether the identified stiffness pattern is both necessary and sufficient for achieving the transition, we performed a sensitivity analysis by perturbing each $\Delta k_{\text{swing}}$ by ±$\pm 5\%$ and observing changes in the objective function. As shown in Fig. 6(B, top), the hindlimb curves (LH and RH) show clear minima at zero perturbation, indicating local optimality. In contrast, the forelimb curves (LF and RF) exhibit much smaller variations. The maximum cost increase was 0.54% for LH and 0.39% for RH, compared to only 0.13% and 0.06% for LF and RF, respectively. These results confirm that the transition is highly sensitive to hindlimb stiffness modulation, while the forelimbs remain largely uninvolved.
To further test the sufficiency of this mechanism, we simulated the reverse transition (RL $\rightarrow$ TR). The resulting stiffness changes were $+12.5\,mgl_{0}$ for LH and $-17.7\,mgl_{0}$ for RH, again showing large, opposing adjustments in the hindlimbs. The forelimb changes remained small: $-4.1\,mgl_{0}$ for LF and $+0.72\,mgl_{0}$ for RF. These results mirror the asymmetry observed in the forward transition and reinforce the conclusion that abrupt hindlimb stiffness modulation is sufficient to switch footfall sequences. The full set of values is provided in Table 3, using the same highlighting scheme.
Together, these results show that the TR $\leftrightarrow$ RL transition can be achieved through hindlimb stiffness modulation alone, without requiring adjustments in the forelimbs. This supports the broader hypothesis that discrete gait transitions in quadrupeds can be driven by targeted, stride-scale mechanical adjustments in a small subset of limbs.
RR $\leftrightarrow$ TR Gait Transition
Figure 7: Replicated gait transition for Individual 1 from rotary right (RR) to transverse right (TR) gallop. (A) Simulated footfall sequence. (B) Swing leg angular velocities. (C) Tugline force trajectory. (D) Absolute cost variation under ±$\pm 5\%$ perturbations in $\Delta k_{\text{swing},i}$. (E) Relative percentage cost variation. These results highlight the dominant role of forelimb stiffness modulation in enabling the transition.
Figure 7(A) illustrates the simulated dynamics for the RR $\rightarrow$ TR transition. The model achieves excellent agreement with experimental footfall and tugline force data, yielding a weighted fit score of $R^{2}_{\text{w}}=0.99$. In the middle panel, the right fore (RF, yellow) initially oscillates faster than the left fore (LF, red), and this asymmetry increases in the second stride—RF becomes even faster, while LF slows. In contrast, the hindlimb angular velocities remain nearly identical across both strides, suggesting that the transition is primarily driven by forelimb stiffness modulation.
This interpretation is supported by the corresponding changes in swing stiffness (Table 3). The forelimbs exhibit strong, opposing changes: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LF}}=+9.5\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RF}}=-15.2\,mgl_{0}$. The hindlimb changes are minimal: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LH}}=-1.8\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RH}}=-0.5\,mgl_{0}$. This pattern, highlighted in the table with shaded forelimb cells, contrasts sharply with the hindlimb-driven transition observed in the TR $\leftrightarrow$ RL case. Sensitivity analysis (Fig. 7(B)) confirms the dominant role of the forelimbs. Perturbations of $\Delta k_{\mathrm{swing},\ i}$ in LF and RF led to cost variations of 0.65% and 0.41%, respectively, compared to just 0.13% and 0.08% for LH and RH. These findings demonstrate that precise modulation of forelimb swing stiffness is sufficient to reproduce the observed footfall sequence transition.
Since the reverse transition (TR $\rightarrow$ RR) was not observed in our dataset, we limited our analysis to the RR $\rightarrow$ TR case. Nonetheless, the consistency of the stiffness modulation and the high reconstruction fidelity reinforce our broader hypothesis that discrete gait transitions can be achieved through stride-scale adjustments to a subset of limb parameters, in this case via targeted forelimb stiffness adaptation.
V Discussion and Conclusion
Gait transitions in high speed galloping
Gait transitions in quadrupeds are typically associated with changes in locomotor speed. High-speed galloping is often viewed as exhibiting consistent, stereotyped stride patterns [4, 22, 5]. In contrast, our analysis of sprint sled dogs reveals frequent, stride-to-stride transitions among distinct galloping gaits that occur without significant changes in speed, stride duration, or terrain.
This behavior was especially prominent in Individual 1, who alternated between gait types under nearly identical locomotor conditions. This observation brings into question the role of multistability in quadrupedal load pulling. Prior studies have suggested that unloaded quadrupeds may switch lead limbs within an asymmetrical gallop to help balance muscular workload and mitigate fatigue [33]. It has also been suggested that sufficiently large perturbations might induce transitions between distinct galloping modes, though empirical efforts have provided inconclusive evidence [6]. In sustained load-pulling tasks, the ability to switch between mechanically distinct gaits may provide sled dogs a functional advantage for maintaining long-duration performance. Within a team, such transitions could also serve to adaptively redistribute load in response to tugline perturbations, contributing to more robust locomotion. However, the consistent gait observed in Individual 2 suggests that such transitions, if they exist, are not strictly necessary.
A minimal model for reproducing stride-to-stride transitions
To understand the mechanisms underlying these transitions, we developed a reduced-order SLIP model with a load-coupled body. Using footfall timing and tugline force data as references, we applied trajectory optimization to replicate all observed galloping gaits and their stride-to-stride transitions. Despite its simplicity, the model successfully reproduced both steady-state gaits and transition dynamics using a small number of adjustable stride-level parameters.
Among the parameters examined, swing leg stiffness emerged as a dominant control variable for enabling gait transitions. Biomechanically, this parameter governs how rapidly a leg swings through the air: higher stiffness corresponds to faster oscillation and earlier touchdown. Our results demonstrate that targeted modulation of swing stiffness in just one limb pair was sufficient to replicate the experimentally observed transitions. Specifically, modulating hindlimb stiffness alone reproduced the transition between TR and RL, while adjusting forelimb stiffness enabled the transition from RR to TR. These adjustments effectively alter the sequencing of leg touchdown events and thus the overall footfall pattern. Sensitivity analyses further confirmed that the cost function was highly responsive to perturbations in the switched limb pair and relatively insensitive to changes in the other limbs. This suggests that sled dogs may exploit stride-wise adjustments in swing timing—achieved through limb-specific stiffness control—as a localized, efficient strategy to switch gaits without global reconfiguration or speed change.
These findings extend previous SLIP-based approaches by showing that transitions between distinct galloping gaits can be replicated without speed variation or terrain change. Instead, stride-scale adjustments to mechanical parameters in a small set of limbs appear sufficient to drive discrete footfall sequence changes, even under external loading.
Limitations and future directions
Although our model replicates footfall timings and tugline dynamics with high accuracy, the underlying causes for the observed transitions remain unresolved. One hypothesis is that transitions are influenced by fluctuations in tugline force. However, across 120 RL strides from Individual 1 and 179 TR strides from Individual 2, most strides—79% and 99% respectively—did not involve a transition. While some transitions coincided with observable force changes, others occurred without any clear variation, as seen in Fig. 6 and Fig. 7. Due to the small number of transitions, statistical evaluation is not currently feasible. Larger datasets with more frequent gait changes and better-resolved force data would help clarify whether load dynamics play a causal role.
One possible explanation is that variations in pulling force influence the energetic interaction between the dog and the sled, thereby prompting gait transitions. Evaluating this hypothesis is limited by the available data. Among 120 rotary-left (RL) strides from Individual 1 and 179 transverse-right (TR) strides from Individual 2, over 79% and 99% of strides remained unchanged, respectively. While changes in tugline force were observed in some transition instances, as shown in Fig. 6 and Fig. 7, similar transitions also occurred with little to no change in force. Again, additional data with more frequent gait switches would be necessary to determine whether load dynamics are a causal factor in gait selection.
Although tri-axial accelerometers were deployed on the same individuals, they were not used in this analysis. Recent work, however, has demonstrated that stance and stride timing can be inferred from inertial measurement unit (IMU) data using machine learning approaches [34], offering a valuable alternative when labeled video is limited or unavailable. To explore this potential, we conducted a preliminary analysis using acceleration data from Individual 1. A bidirectional LSTM model [35] trained on Individual 1’s labeled dataset was able to identify footfall sequences and stance timing with high precision. Across 362 footfalls in a held-out test set, stance events were detected in the correct order, with an average intersection-over-union of 0.869 between predicted and ground truth stance intervals (corresponding to an average timing error of $\sim 10\text{ms}$). A full description of this analysis, including the Python model, training procedure, and evaluation, is provided in . While promising, the temporal precision was insufficient for use in our model fitting. The limited training dataset also precludes broader generalization. Nonetheless, this approach demonstrates a promising path toward accurate, field-deployable gait classification in scenarios where continuous video capture is impractical.
Another limitation of this study lies in the inability to decouple the roles and contributions of individual dogs within the larger team. All experimental data presented here were collected from a ten-dog team hitched in a double-tandem (Alaska) configuration, where the distribution of tensions across the interconnected tether network is not directly measurable. This introduces some ambiguity in interpreting how load dynamics influence gait selection at the individual level. In practice, each dog’s pulling contribution may vary depending on spatial position, terrain conditions, cornering behavior, fatigue, or even team coordination strategies. Whether such factors contribute to observed gait transitions remains an open question.
Future studies could address this limitation by instrumenting each segment of line to measure complete hitch network tension in real time. Additionally, synchronized video and acceleration data from multiple team members could enable modeling of team-level coordination strategies, including how leading versus trailing dogs modulate their gait or force output in response to group dynamics. Decoupling these roles would allow for more precise attribution of observed gait transitions to local mechanical or environmental factors, and could clarify whether transitions are driven by individual control decisions or emergent team-level phenomena.
Acknowledgements.
We gratefully acknowledge the generous support of the sled dog mushers who volunteered their kennel and dogs for data collection. Their willingness to share their time, expertise, and home made this fieldwork possible. Funding: S.B. acknowledges funding from NSF CAREER IOS-1941933 and Schmidt Sciences, LLC. Author contributions: Conceptualization: J.D., B.S. Experiments and related data analysis: B.S. Model and related analysis: J.D. Writing—original draft: J.D., B.S. Writing—review and editing: J.D., B.S., S.B., Z.G. Visualization: J.D., B.S. Supervision, funding acquisition: S.B., Z.G. Competing interests: The authors declare no competing interests. Data and materials availability: All code to run simulations, generate figures, and reproduce results is publicly available at ….